OpenStudy (anonymous):

The set of real numbers between 0 t0 1 should be finite or infinte and explain it by reason

OpenStudy (kc_kennylau):

infinite, and uncountable as well

OpenStudy (anonymous):

how please explain me,

OpenStudy (kc_kennylau):

First let's look at the definition of countability: "A set S is called countable if there exists an injective function f from S to the natural numbers N = {0, 1, 2, 3, ...}." (Copied from Wikipedia)

OpenStudy (anonymous):

ok then

OpenStudy (kc_kennylau):

Let's suppose that the set is countable.

OpenStudy (kc_kennylau):

And define a function f(x) such that for every natural number x, f(x) refers to some real number between 0 and 1

OpenStudy (anonymous):

@kc_kennylau i just started sets and i just find the definition of it just tell me that o and 1 between these two what the numbers are present

OpenStudy (kc_kennylau):

For example, \[ f(0)=0.\color{red}3363684485...\\ f(1)=0.6\color{red}955102578...\\ f(2)=0.26\color{red}39483653...\\ f(3)=0.006\color{red}5093130...\\ f(4)=0.8864\color{red}533021...\\ f(5)=0.84447\color{red}42908...\\ f(6)=0.862720\color{red}1508...\\ f(7)=0.8665645\color{red}627...\\ f(8)=0.39569269\color{red}52...\\ f(9)=0.824607139\color{red}4...\\ \] Now, let's take the first digit of the first number, second digit of the second number, third digit of the third number, etc... and form a new number: \[0.3935541654...\]

OpenStudy (kc_kennylau):

Now add one to all digits of this (if it's 9, convert it to 0) \[0.4046652765...\] @ziqbal103 my modem was down :/

OpenStudy (anonymous):

ziqbal, your question was just about the set being finite or infite. Don't get stuck on the countability therefore. :) The set is inifite as already the rational numbers betwee 0 and 1 are infinite. Which you can prove very nicely by just writing them all down, each divisor in one line.

OpenStudy (kc_kennylau):

@techhelper @ziqbal103 oh i see sorry

OpenStudy (kc_kennylau):

To prove that the set of all real numbers between 0 and 1 is infinite: The first digit has 10 probabilities. The second digit has 10 probabilities. The third digit has 10 probabilities. And so on and so on. Therefore, the number of all real numbers between 0 and 1 =10*10*10*..... =\(\infty\)

OpenStudy (anonymous):

@techhelper tell me how the rational numbers become infinte

OpenStudy (anonymous):

and @kc_kennylau the difference b/w o and 1 there is more than 10 decimal digits i means infinite so am i right

OpenStudy (kc_kennylau):

@ziqbal103 sorry I fail to comprehend your last sentence.

OpenStudy (anonymous):

Assuming they are finite, you can write them in a sequence that is always getting bigger: 0, a1, a2, a3, ... , 1. take any pair of them like a2 and a3 and have a look at a number between them: b=(a2+a3)/2. this is by definition a rational number again. however you just found a new one, that is not yet in your sequence, so your assumption of them being finite must be wrong.

OpenStudy (kc_kennylau):

@techhelper nicely done :D

OpenStudy (kc_kennylau):

For your real number question: Assuming they are finite, you can write them in a sequence that is always getting bigger: 0, a1, a2, a3, ... , 1. take any pair of them like a2 and a3 and have a look at a number between them: b=(a2+a3)/2. this is by definition a real number again. however you just found a new one, that is not yet in your sequence, so your assumption of them being finite must be wrong.

OpenStudy (anonymous):

@techhelper and @kc_kennylau now i get it thank you very much for helping

OpenStudy (kc_kennylau):

no problem :)

OpenStudy (anonymous):

i wanna ask another question

OpenStudy (anonymous):



OpenStudy (kc_kennylau):

\[\large z_1z_2=(a+ib)(c+id)\\\large\overline{z_1z_2}=(a-ib)(c-id)\\\large z_1z_2\overline{z_1z_2}\\\large=(a+ib)(a-ib)(c+id)(c-id)\\\large=(a^2+b^2)(c^2+d^2)\\\large=|z_1z_2|^2\]

OpenStudy (kc_kennylau):

I mean \(\large|z_1|^2|z_2|^2\)

OpenStudy (kc_kennylau):

@Directrix @ganeshie8 just wanna check if I'm correct thanks

OpenStudy (anonymous):

i think same case will be with(z1+z2)(z1.z2whole bar)....?

OpenStudy (anonymous):



OpenStudy (kc_kennylau):

Sorry I'm not familiarized with complex number :/

OpenStudy (anonymous):

ok@kc-kennylau.and @techhelper you?

OpenStudy (anonymous):

@ehuman

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (anonymous):

@kc_kennylau where are you help me i am sending you a link an attachement so that you make me clear that z1+z2whole modulus is less than z1modulus +z2 modulus ATTACHEMENT

OpenStudy (anonymous):

http://www.suitcaseofdreams.net/properties_modulus.htm BY THIS LINK CAN YOU HELP ME @kc_kennylau

OpenStudy (kc_kennylau):

Sorry I'm not familiarized with complex number

OpenStudy (kainui):

This is my reasoning: Take a number between 0 and 1. Divide it by 2. Divide the next number by 2. Keep dividing until you're convinced you will keep getting numbers no matter how much you divide.